Find any inflection points and the intervals of concavity for the function y = 3x-1 x2
Answer #1

and concavity : where occurs Inflection points Inflection second derivative concarita Changes. points j.e.) where the zero 2=-36-2) 3-2-1 E =36) 3 (-2) x-2-1 +2x-3 >> 4- =-30-2 +2x-3 y d (Y) + 2x-3 2 d da Р P 34-2+2x+3] (- +2 6 6 .4 -3 6 x -6.3 X4x=1 is a point, because inflection concarity Changes evaluations from above Hence 30-1 2 32-1 = y(i)= Y- x2 2 cr (1) Inflecti